OpenStudy (anonymous):

16n^2-1=0 n=? solve and check help please

jimthompson5910 (jim_thompson5910):

Hint: Use the difference of squares factoring rule a^2 - b^2 = (a-b)(a+b)

OpenStudy (anonymous):

(4n+1)(4n-1)??

jimthompson5910 (jim_thompson5910):

good, so 16n^2-1=0 turns into (4n+1)(4n-1) = 0

jimthompson5910 (jim_thompson5910):

Now use the zero product property to break up that last equation to get 4n+1 = 0 or 4n-1 = 0

OpenStudy (anonymous):

So the answer would be -4/1,4/1

jimthompson5910 (jim_thompson5910):

close, but no

jimthompson5910 (jim_thompson5910):

4n+1 = 0 4n =-1 n = -1/4 --------------- 4n -1 = 0 4n = 1 n = 1/4 So the solutions are n = -1/4, n = 1/4

OpenStudy (anonymous):

cool thank you that makes sense thanks alot for breaking it down

jimthompson5910 (jim_thompson5910):

np